An electric motor uses 180m of copper conductor with a cross sectional area of 0.26 cm². Find the resistance of the electric motor. The resistivity of copper is 1.72x10^-8 ohm meter.


1
Expert's answer
2021-03-25T19:28:40-0400

We can find the resistance as follows:


"R=\dfrac{\rho L}{A},""R=\dfrac{1.72\cdot10^{-8}\ \Omega\cdot m\cdot10\ m}{0.26\ cm^2\cdot(\dfrac{1\ m}{100\ cm})^2}=0.12\ \Omega."